A particle moves along the parabolic path x=y2 + 2y + 2 in such a way that the y component of velocity vector remains 5m/s during the motion. The magnitude of the acceleration of the particle is? 

Grade:11

1 Answers

Khimraj
5 years ago
x = y2 + 2y +2
differentiating wrt time
vx = 2yvy + 2vy
again differentiating wrt time
ax= 2vy2  + 2yay + 2ay
ay = 0 as vy = 5m/s(constant)
then ax = 2*52 = 50 m/s2
So a = \sqrt{a_{x}^{2} + a_{y}^{2}} = 50m/s2
Hope it clears.